OpenStudy (anonymous):

whats the value of x in the following equation? (x + 4) + 3 = 2x( + )

OpenStudy (anonymous):

not this one srry this is the equation

OpenStudy (anonymous):

u missing sumthin in the RHS?

OpenStudy (anonymous):

x+7=2x

OpenStudy (anonymous):

1/2(x+4)+3=2x(9/2+3/4)

OpenStudy (anonymous):

that is the real equation

OpenStudy (anonymous):

x=1/2

OpenStudy (anonymous):

how bout this one?

OpenStudy (anonymous):

14x − 6 = −3 + 2x + 5

OpenStudy (anonymous):

x=2/3

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

and this word equation

OpenStudy (anonymous):

The instructions for baking a turkey recommend cooking a 12-pound bird for 30 minutes in a 350 degree oven. The instructions also suggest adding 3 extra minutes of cooking time for each additional pound. The cooking time for a turkey can be calculated using the equation below, where w represents the wieght of the turkey. If Maria needs to cook her turkey for 45 minutes, how many pounds does Maria's turkey weigh? Cooking Time = 3(w - 12) + 30

OpenStudy (anonymous):

hello?